OpenStudy (anonymous):

solve for x. a) 27^5x+4=9^x-12 b) 4^x+2=24

OpenStudy (anonymous):

i would try and change to the same base first

OpenStudy (anonymous):

a) 3^3(5x+4) = 3^2(x-12) so 3(5x+4) =2(x-12) etc

OpenStudy (anonymous):

wow, nice job mathwiz you can defnintly solve these kinds of problem

OpenStudy (anonymous):

in the 2n one 4^(x+2) = 24?? (x+2) ln 4 = ln(24) x+ 2 = ln(24)/ln4 x= -2 + ln(24)/ln(4)

OpenStudy (anonymous):

unless u made a typo ... can't convert the 24